R is the midpoint of PS.Q is the midpoint of PR. PR=3. what is the length of QS
1
Expert's answer
2020-11-01T17:26:02-0500




"R" is the midpoint of "PS" .

"Q" is the midpoint of "PR." .

"PR=RS=3"

"QR=\frac{1}{2}PR"

So,

"QR=\frac{1}{2}×3=1.5"


"PR=RS"

therefore,

"PR=RS=3"

"QS=QR+RS"

"QS=1.5+3=4.5"

"QS=4.5(Answer)"